Commissioning Notes from 14-15 May 2008 MST

Observers: J. Hill, J. Storm and N. Cushing on the mountain, plus D. Miller and A. Rakich from Tucson
Telescope Operator: R. Grashuis

Summary

Scattered rain showers and scattered TCS problems gave the night a very slow start. I updated the pointing and collimation models to correct the consistent difference in X that we see compared to last month. Clear sky and subarcsec seeing at 2AM turned it into a quite productive night. We took more data to calibrate the guide probe position. We were able to guide successfully on off-axis stars at several position angles. However, guiding gets wobbly sometimes. We think this is because we are not correcting for the probe rotation angle for these stars. We learned that the spots are vignetted on one side of the WFS pupil at the maximum field angle. The rotator quit working at 09:23, and we couldn't revive it from the MCSPU interface. Doug sent the first IDL wavefront correction for an on-axis star about 09:40 UT, but the zernikes were wrong because he was reading the rotator demand from DD rather than the actual stuck rotator angle. We spent about an hour working on the S-H correction of astigmatism and focus. We took some images at the end of the night in NGC6341 to provide astrometric calibration of the guider field in decent seeing.

Nighttime Activities

01:15 Open Dome for equilibration. T=6degC, Wind light. Sky filled with heavy cirrus, but it is not snowing.

02:45 pointto M5_0441, but the preset timed out.

Doug reports that he can't take CCD images from AGW.

Randy had problems with MCS on lbtmu101, but it works on lbtmu102

GCSL is crashing immediately after starting it.

03:05 stopAGW, startAGW didn't help with taking images.

03:12 stopAGW again, stop oacserver and check4nports

03:13 reboot agw-control

03:16 startAGW after oacserver and check4nport, and now Doug can take images.

03:20 GCS command line commands still aren't working (i.e. getStatus, selectAGW). Somewhere in here Michele restarted network services and GCSL on lbtmu107. That seems to have GCS running again.

Send Z8=+2000 to PSFL for an initial guess on collimation.

03:29 start spiral search, but now IRTC doesn't seem to be adjusting the exposure time???? or it isn't giving us the new image. An apparently 6 sec H exposure is still not saturated.

Close

03:46 Close dome because of an isolated tiny cloud spitting rain just over us.

03:47 reboot the irtc windows computer. That seems to have cured the IRTC problems. Although I'm not sure exactly what the problems were. It seemed like the camera was not listening to commands from the GUI (or vice versa).

03:58 preset to M5_0619 gstar=5 GUIDE to test today's new GCS. GCS now has the correct probe position of 33.3 506.4. Still 20 minutes later this command had not finished. So we had to kill and restart irs.

04:25 We loaded an LBC pointing model. And we note that you can send "ptmodify IE -28" or "ptmodify IE -10" to this different pointing model. So the numerical bug is pointing model dependent.

04:4x Move OSS, MCS, PSFL back to lbtmu101 since Chris and Michele can't see anything wrong with 101. [And they ran there without problems the rest of the night.]

Open

05:18 Rain is gone, although we still have 3 mags of cirrus and lots of moon. Reopen since these commissioning tests can work on fairly bright stars.

05:26 pointto M5_0619 for initial collimation. This preset hangs waiting for something. kill irs at 05:36.

05:37 Randy presets telescope with presetGUI and that works.

05:52 Stop and Restart CSQ on lbtmu103. That didn't help.

Called Chris Biddick

05:58 CSQ was hung waiting for GCSL on the stop guiding command. GCSL was present but not responsive. Restart GCSL and we can preset again.

Preset to M5_0619, used spiral search to hunt for star. Found at offset -60 30.

Telescope is collimated at nearly the same position we have had for the last few clear nights:
  • adjustPointing left M1 -15 0
  • Z8=+3500 Z7=200.
  • Offset zeroed with CA=-108 IE=-38

06:21 pointto M5_0619 gstar=1 PA=0 GUIDE, Guide acquisition fails because of heavy clouds and because of the default 1 msec exposure.

Close

06:24 Close Dome for heavy clouds and threat of rain.

New SX Pointing and Collimation Models

We have seen consistent and large displacements in X/RY over the last couple nights of observing compared to last month. Since these collimation results have been consistent, I have added them into the pointing and collimation models.

Installed the new pointing model irtc.20080514.ptmod (which is irtc.20080415.ptmod with CA=-108 and IE=-38) which is in /home/telescope/TCS/Configuration/PCS.

And I have calculated new global offsets for the collimation model based on the sum of our adjust pointing corrections (which provide additional range in +X) and the active optics corrections to correct the large amount of X-coma that we see.

Active X=+2.988 Y=+0.171 Z=0.0 RX=1.8 RY=-32.1
Point X=-0.698 Y=0 Z=0 RX=0 RY=15
New Global X=+2.29 Y=0.17 Z=0 RX=1.8 RY=-17.1

/Repository disk cleanup

Norm and John are cleaning out space on the /Repository disk which was 93% full. The biggest nights of LBC observing are 32 GB. We've deleted the Dec07, Jan08, and Feb08 LBC data from /Repository/Check_and_Delete. /Repository is down to 68% full.

Open

08:05 Open again in a rapidly clearing sky

IRTC is at 240 K and seems to run happily there.

M5_0619

08:15 guidiing was jumpy when the exposure was long. Saturation? Speckles? But, we were guiding off axis!

08:24 stopGuiding

08:25 pointto TRACK (in order to recenter the telescope and leave the guide probe deployed)

guider000106 star=1 xy -12.48 541.177

Note that all these images need to be corrected for image location relative to the hotspot before refitting the probe coordinates. Also note that we were able to find almost all the guide stars with the probe using the first rough transformation.

08:xx pointto gstar=2 GUIDE

08:30 vibration of image

08:34 Study this guider sequence. After cogitating on it for a while, we suspect that guiding is jumpy when the probe is away from x=0, and GCS/PCS is not correcting for the probe angle.

pointto TRACK

guider000108 star=2 xy 65.88 567.12

08:35 pointto gstar=5 GUIDE

pointto TRACK

guider000109 xy 33.59 506.53

08:39 pointto gstar=2 PA=90 GUIDE Guiding did not run awy, but is surely not making the proper correction here at PA=90, and X=-45.73.

pointto TRACK

guider000110 xy -45.73 546.54

08:43 pointto gstar=8 PA=90 GUIDE GCSL died unexpectedly.

08:45 had to give "netconfig stop GCSL" in order to start GCSL that we just restarted.

guider000111 xy 44.75 451.79

08:48 GCSL started properly

08:49 pointto gstar=10 PA=90 GUIDE, Didn't find star = clouds?

pointto TRACK

guider000112 xy 104.08 462.83

08:53 pointto gstar=11 PA=180
May 15 01:52:53 lbtmu107 LBT_left_GCS: converted into native AGW coordinates by GCS (x,y) = -121.063773, 461.256672 mm
May 15 01:52:56 lbtmu107 LBT_left_GCS: error: setProbePosition: couldn't move probe to position (SFP) (x,y): -121.063773, -151.243328 mm.

08:55 pointto gstar=6 PA=180 GUIDE - couldn't find star

guider000114 xy -5.09 457.39

08:58 pointto gstar=7 PA=-90 GUIDE, guides OK here although lots of atmospheric speckle motion. Doug and Andrew got a wavefront from this star.

Star ran away on the "TRACK" preset. Had to send it again.

guider000115 xy 30.92 451.41

09:04 pointto gstar=12 PA=-90 GUIDE, couldn't find star (too near edge)

guider000116 xy 2.42 420.42

09:07 pointto gstar=13 PA=-90 GUIDE (guiding and WFS here) Seeing has deteriorated

09:09 Presetting to pointto "TRACK" while guiding stops guide thread, and the telescope goes to unknown far away place (outside the 30 arcsec field).
May 15 02:09:14 lbtmu107 LBT_left_GCS: guide thread stopped. GCSL did not report seeing the stop guiding command from CSQ.

09:10 pointto TRACK

guider000117 xy -56.98 422.87

09:13 pointto gstar=9 PA=0 GUIDE
May 15 02:15:59 lbtmu107 LBT_left_GCS: fatal error: did not receive full CCD readout from guide camera within timeout period of 3659 ms.

guider000119 xy -55.66 444.626 500ms

setxy -x 0.0 -y 612.5 call failed while checking bits on UMAC: RPC: Timed out

getfilter oac=3 (blank)

guider000121 still saturated at 1 ms

setfilter -f 0

guider000122 1 ms filter oac=0

M5_0738

09:23 pointto M5_0738 gstar=4 PA=180 GUIDE (rotator didn't slew to track???? why won't it move?)

Randy uses preset GUI to see if he can run rotator. It won't work. Something about brakes timing out. Issuetrak 1351.

BS9177

09:38 pointto BS9177 gstar=0 PA=0 GUIDE (probably we have to wait for the preset to time out, no it fails for no rotator motion)

Center star on IRTC with CA=-98.5 IE=-40.5

guider000124? (image at 296,255) based on centering on IRTC.

guider000126 result of first wavefront correction is 3 arcsec diameter. So needs work on the corrections.

recenter with IRTC CA=-97.5 IE=-41.0

guider000127 (297,245)

10:04 Focus has been 21, now set to 26

guider000129 foc=26 image is better

guider000130 foc=31 image maybe slightly better

guider000131 foc=36 image maybe slighty worse

guider000132 foc=41 image similar to 21 with astigmatism rotated

guider000133 Adopted focus =31 based on John's eyeball of the image sequence, star at (301,256) (not best focus for IRTC)

10:11 Taking more S-H WFS data.

10:14 Doug sent -544 Z4

guider000134 (308,254)

Doug sent -643

guider000135 (better)

10:24 recentered with CA=-95

./readGuideCam -e 100 (some 1/4 of these commands hang and you have to ^C out of them, and on readWFS) At least one time it said: Image Server: Could not bind to port 10.144.0.64:30000

10:30 Doug sent Z4=+68.5 nm

Doug sent Z5=566 Z6=-232

10:45 Discover that Doug is reading rotator demand from DD, and not the actual position of the stuck rotator.

10:54 IE=-43 CA=-95.5

11:09 guider000145 - they are making progress on correcting focus and astigmatism. Measure this one.

11:26 guider000146

11:28 preset to NCG63xx star G

guider000147 5 sec B

guider000148 10 sec B

setfilter -f 2 (r')

guider000149 15sec r'

guider000150 30 sec r' (demand 64.795 for PA=0, compared to 99:08:08 deg where rotator is stuck.)

guider000151 30 sec r'

guider000152 30 sec r'

dither IE=-45 in elevation

guider000153 30 sec r'

dither IE=-38 in elevation

guider000154 30 sec r'

guider000155 30 sec r'

dither CA=-90

guider000156 (demand 66.198) 30 sec r'

dither CA=-100

guider000157 30 sec r'

dither CA=-110

guider000158 30 sec r'

11:48 point at core of cluster AND refocus!!!!

guider000159 probe still moving

guider000161 10sec r'

Clear Active Optics and focus quickly by eye on IRTC

setfilter -f 0 (B)

guider000166 5 sec B

guider000165 5 sec B plus bonus sky flat

11:57 trip MCR relay while moving building in azimuth to close. Sky is mostly clear with just a few scattered clouds.

Copied guider0001??.fits images from /tmp on lbtmu107 to /Repository/AGW_Data/080515g/

Plan for tomorrow night

.Tonight...Partly cloudy. Chance of rain showers with chance of thunderstorms then chance of rain showers. Probability of measurable precipitation 40 percent. Northeast wind 12 mph. Low 35.

Jesper is going home, and with a forecast like this John is probably going home too.

Plan for the next night that is clear.

Doug and Andrew will continue working on the convergence of the active optics corrections.

-- JohnHill - 15 May 2008
Topic revision: r4 - 15 May 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback